0 0
Add a comment Improve this question Transcribed image text
Answer #1

a(t)=cos t-sin t

dv/dt=cos t-sin t

integrating,
v(t)=sin t+cos t +c

where c is a constant

now a t=0,
v=1

so c=0

v(t)=sin t+cos t

dx/dt=sin t+cos t

x(t)=-cos t+sin t+c


x(0)=0

so c-1=0

c=1

hence x(t)=sin t-cos t+1


b)v(t)=sin t+cos t

v(t)=0

==> sin t=-cos t

tan t=-1

t=-pi/4 or 3*pi/4

c)x(t)=sin t-cos t+1

plotting it for t=[0,5] we can see that

minimum occurs t at=5 and maximum occurs at dx/dt=0

i.e. t=3*pi/4=2.356 seconds

minimum value=-0.24259

maximum value=2.4142